OpenStudy (marsmcnugget):

Which expressions show how 6 • 45 can be rewritten using the distributive property? Choose all answers that are correct. A. 6 • 40 + 6 • 5 B. 6 • 40 − 6 • 5 C. 6 • 50 – 6 • 5 D. 6 • 50 + 5

OpenStudy (anonymous):

6*5= 30 6*40+6.5= 270 the answer 6*50-6*5=210 not the answer 6*50-6*5=270 the answer 6*50+5=305 not the answer